OpenStudy (anonymous):

What is the completely factored form of t4 − 16?

pooja195 (pooja195):

ok did u try it?

OpenStudy (anonymous):

Yes. I got (t2-4)(t2+4)

pooja195 (pooja195):

Thats correct :)

OpenStudy (anonymous):

ok thanks:)

OpenStudy (anonymous):

Wait! No it's not. I got it wrong:(

pooja195 (pooja195):

(t2+4)(t+2)(t−2)

pooja195 (pooja195):

i was thinking both of them are the same thing

OpenStudy (anonymous):

Ohhh okay :)

pooja195 (pooja195):

what was the right answer?

OpenStudy (anonymous):

Yes:)

pooja195 (pooja195):

:) welcome to openstudy :)

OpenStudy (anonymous):

haha thanks :)

OpenStudy (mathstudent55):

Factoring means factoring completely. When you have to factor a polynomial of two terms, look for the difference of squares or the sum or difference of cubes. The first step in factoring, though, is always to look for a common factor. In this case there is no common factor. Then you have two terms. t^4 is obviously not a cube, so look at the difference of two squares. Sure enough, \(t^4 - 16\) can be written as \( (t^2) - 4^2 \) which clearly looks like the difference of two squares.

OpenStudy (mathstudent55):

The difference of squares factors like this: \(a^2 - b^2 = (a + b)(a - b) \) Your expression factors similarly: \((t^2)^2 - 4^2 = (t^2 + 4)(t^2 - 4) \) Now you need to look at each factor to see if it can be factored further. \(t^2 + 4\) is a sum of squares. It can't be factored. Now we look at \(t^2 - 4\) which is a difference of squares and can be factored. \(t^4 - 16 = (t^2)^2 - 4^2 = (t^2 + 4)(t^2 - 4) = (t^2 + 4)(t + 2)(t - 2) \) Finally we look at the new factors, t + 2 and t - 2, and we see that they can;t be factored any further, so our answer above is fully factored.